# Meridional Circulation Streamlined

### v1

@@ Appendix B Vertically varying radiative diffusion
We use Dedalus2 to solve equation (15) with $k$ varying in the vertical, which is the more generally relevant scenario. We see the steady-state zonal wind profiles in Figure 6 (right panel) corresponding to different depths at which radiative diffusion is shut off (left panel). That radiative diffusion is the clear driver of deep penetration is evident from the truncation of the spread of zonal wind in the lower region for models with larger Prandlt number. Another feature is the accumulation of kinetic energy in the strongly radiative-diffusive upper region, leading to higher-magnitude winds. The shallower this region of energy accumulation, the stronger the wind.

### v2

@@ Appendix B Vertically varying radiative diffusion
We use Dedalus2 to solve equation (16) with $k$ varying in the vertical, which is the more generally relevant scenario. We see the steady-state zonal wind profiles in Figure 5 (right panel) corresponding to different depths at which radiative diffusion is shut off (left panel). The $k$ profiles are chosen to be nearly non-differentiable so that maximum effect on the solution may be observed. That radiative diffusion is the clear driver of deep penetration is evident from the truncation of the spread of zonal wind in the lower region for models with larger Prandtl number. Another feature is the accumulation of kinetic energy in the strongly radiative-diffusive upper region, leading to higher-magnitude winds. The shallower this region of energy accumulation, the stronger the wind.